OpenStudy (anonymous):

39. Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 2 = 1/3 (x – 5); (–4, 3) (1 point) y = –3x + 15 y = –3x – 9 y = –4x + 3 y = –1/3 x – 11

OpenStudy (anonymous):

It's the second one because you take the negative reciprocal of the given slope to get a new slope: -3 Now, you use the point-slope formula: y - y1 = m(x - x1) y - 3 = (-3)[x - (-4)] y = –3x – 9

OpenStudy (anonymous):

Just... Thank you, I truly do appreciate it.

OpenStudy (anonymous):

You're always very welcome and it's a great privilege working with such a nice person as yourself! @sgstiles